OpenStudy (anonymous):

Suppose I had to find two light bulbs from a collection of 12, all but 3 of which are working. If I test each one in turn, what is the probability that I would find two working bulbs (a) among the first three bulbs tested (b) when three bulbs have been tested, but not before?

OpenStudy (anonymous):

probability that any one bulb is working is \[\frac{9}{12}=\frac{3}{4}\] and probability it is not working is \[\frac{1}{4}\] if you pick three at random the probability that exactly two are working is \[3(\frac{3}{4})^2\times(\frac{1}{4})\]

OpenStudy (anonymous):

bro, the bulb tht is working is 3/12 not 9/12 read the question :)

OpenStudy (anonymous):

hmmm how does a bro interpret the statement "ALL BUT THREE OF WHICH ARE WORKING"?

OpenStudy (anonymous):

wch means 9 are working out of 12. are you sure bro?

OpenStudy (anonymous):

How to do the (b) part? Can you please explain?

OpenStudy (anonymous):

satellite is wrong

OpenStudy (anonymous):

So what's the correct answer?

OpenStudy (anonymous):

I think satellite is right for the statement all but 3 is working means 3 are not working